# active distribution network DER distributed energy resource

### v2

@@ Abstract
Index Terms:
Flexibility Aggregation, Ancillary Services, Optimal Power Flow, Robust Optimization
@@ I-B Flexibility aggregation approaches
We consider a DSO-TSO cooperation where the DSO aggregates the DER capabilities to offer flexibility to the TSO. The relevant problem is to quantify and explicitly determine the available flexibility, while accounting for network and resource constraints and stochastic boundary conditions.  [^1] In this respect, the considered problem is different than approaches presented in [8] and [9], where the operation of a distribution system under uncertainty is considered and works such as [10], where clearing of energy and flexibility markets is considered while accounting for DERs. In this section, existing approaches for flexibility aggregation are discussed and the remaining issues are identified. A first distinction between the methods reported in literature can be made based on the type of algorithm used [11]. The first group consists of approaches based on Monte Carlo (MC) simulations. Here, a significant number of scenarios are considered, each corresponding to a feasible actuation of the flexible resources. For each scenario, the corresponding power exchange at the GCP is obtained. If all constraints are satisfied, the corresponding power exchange is labeled as a feasible point within the PQ plane. By performing a large number of simulations, an accurate mapping of the power capability can be made. An example of such an approach is presented in [12]. However, as expected, the authors report that many scenarios are needed to guarantee an accurate representation of the flexibility set, leading to large computation times, and that additionally it is hard to capture the extreme points of the capability map. Recent developments (e.g. [13]) consider this problem and propose a Bayesian learning approach to alleviate the computational challenges. However, they still rely on an initial outer-box approximation of the flexibility set obtained through optimization.
@@
On the other hand, optimization-based approaches can be divided in a group of methods that attempt to compute the exact power aggregation set at the GCP and a group of methods that compute an approximate set, which can typically be described more easily. We further differentiate existing methods based on four important features of the flexibility aggregation sets: grid-awareness, time-coupling, cost-awareness and wether they account for uncertainty. As an example, [14] proposes an approach to compute the exact aggregated power flexibility set of multiple DERs without accounting for grid constraints. They also show that the number of constraints may be too large to solve the exact problem and have instead proposed a set of approximate models. Recognizing this issue, [15] proposes to cluster DERs based on the similarity of their capabilities and obtain an inner approximation through a specific type of polytope. The DERs in each cluster can then be exactly aggregated. Alternatively, in [16] a ”geometric prototype” is constructed by selecting a subset of the equations forming the exact power aggregation set, keeping geometrical features of the exact model while reducing the computational complexity. Approaches attempting to identify the exact aggregated flexibility set, while simultaneously satisfying the grid constraints, also exist. In this case, a set of OPF problems is typically solved. For instance, in [17] a set of non-approximated non-convex OPF problems are solved to obtain points on the aggregated capability curve in the PQ plane until a convergence criterion based on the distance between the obtained points is satisfied. This approach also allows to include cost constraints. Similar approaches are also proposed in [18], [19], [20] and [21]. Explicit computation of flexibility costs for these approaches is proposed in [22], but costs are computed a posteriori for identified points of the flexibility map and consider a single time step. These approaches only allow to compute the aggregation set for a single time step as the solved OPF problems cannot account for time coupling. To alleviate this problem, [23] proposes a network-informed data driven approach based on a classifier labeling samples to obtain an approximation allowing for temporal coupling.
@@
While the above methods can obtain flexibility areas closer to the exact set, they are only applicable when neglecting grid constraints or considering a single time step or have to resort to approximations to make the problem tractable. Therefore, methods seeking convex inner approximations of the exact flexibility set are desirable. An approach for the estimation of a robust aggregation set is presented in [24], where a linear power flow model is used. However, only single time steps are considered and DERs costs are not taken into account. In [25] a grid-aware approach, using a fixed-point linearization, is presented to compute the aggregate power flexibility over multiple time steps by approximating it as a hyperbox. However, uncertainty of stochastic prosumption is not accounted for. The authors attempt to model the costs, but only the one of the base trajectory, around which flexibility is offered, is included. In [26] the authors show that the flexibility aggregation set obtained through a hyperbox approximation is very conservative and that a larger one can be obtained with an inner ellipsoidal approximation. Using the same fixed-point linearization, the authors present tractable reformulations to obtain the maximum-volume ellipsoid under affine and quadratic disaggregation policies, while accounting for uncertainty through ellipsoidal uncertainty sets at each time step. An approximate approach decomposing the flexibility aggregation set in a ”virtual battery” and ”virtual generator” is presented in [27]. Uncertainty is accounted for through chance constraints assuming Gaussian probability distributions suitably transformed to deterministic constraints. The authors claim better results compared to the ellipsoidal inner approximation. The same authors extend this in [28] with a piecewise fitting of the aggregated cost function for the different time steps. However, the cost function is computed by minimizing the operational cost for different points on the time-decoupled aggregation sets and fitting an affine function. Therefore, it does not account for the time-coupling of the flexibility and its influence on the costs. An alternative approach to the approximate aggregation problem is proposed in [29]. By solving the disaggregation problem for samples selected in the flexibility aggregation space and labeling feasible points a convex ellipsoidal classifier is found. This ellipsoid is used as a surrogate solution space within which a polytope mapping the aggregated flexibility is determined. Uncertainty is accounted for when constructing the surrogate set. All the works above consider the flexibility aggregation problem to be the mapping of a single flexibility commodity at the ADN GCP. However, even though all flexibility services require power exchange, the provision of multiple ones (e.g. primary frequency regulation (FCR) and aFRR) have different implication on time-coupling and, therefore, the feasibility of an aggregation set. This is demonstrated in works regarding the optimal provision of AS using battery energy storage systems, such as [30]. Furthermore, in the works considering uncertainty, forecast errors are either modeled robustly, leading to overly conservative solutions, or through chance constraints related to individual time steps. This does not accurately represent the impact of stochastic variables. Finally, although some works consider costs, none of these methodologies enforce cost-effectiveness of flexibility in an appropriately time-coupled manner. To summarize, none of the above approaches consider simultaneously network constraints, time-coupling and cost of the flexibility and the impact of uncertainty. The provision of multiple AS is also not considered in state-of-the-art aggregation approaches. Therefore this paper makes the following contributions:
@@
    Uncertainty is accounted for in a probabilistic way through joint ellipsoidal uncertainty, following the P90 requirement [31] for the availability of power flexibility.
@@ II Problem Statement
The problem considered in this work is the flexibility aggregation at the GCP of DERs within ADNs with generic topologies (i.e. either meshed or radial) to provide multiple AS at the transmission level. To accurately model the available flexibility, it is essential to consider both the capabilities of DERs and the constraints of the network. This ensures the aggregated flexibility does not adversely impact the operation of the distribution system. Additionally, the impact of uncertain prosumption and the time-coupling of the DERs flexibility capabilities is highly relevant. Consider the general formulation (1), where $p_{0}$ is the vector representing the power exchanged at the GCP over all time steps, $\Omega_{p_{0}}$ is the set of feasible $p_{0}$ values, $p$ and $q$ are stacked vectors containing the vectors of controllable active and reactive power injections [^2] Note that in this formulation, the controllable injections are assumed to be known, with a capability curve that is constrained by the uncertainties. over all time steps and $\zeta$ is the stacked vector collecting the uncertainty drivers $\zeta_{t}\in\mathcal{U}_{t}$ over all time steps.
$\max_{\Omega_{p_{0}},p,q}$ $\quad J(\Omega_{p_{0}}):\forall\zeta_{t}\in\mathcal{U}_{t},t=1..T$ (1a)
s.t. $G(p,q)+b(\zeta)=p_{0},\quad L(p,q)\leq z(\zeta)$ (1b)
$J$ represents the flexibility maximizing objective, $G$ and $b$ represent the mapping between respectively the controllable resources and the uncertainty drivers and the slack power, while $L$ and $z$ represent the network and resource constraints. The goal of (1) is to determine the flexibility-maximizing set $\Omega_{p_{0}}$ of power exchanges at the GCP. Any point within this set should be feasible for any realization of the uncertainty drivers $\zeta_{t}$. Directly obtaining such a set is in general untractable, even when neglecting grid constraints and using linear DER models [32]. The authors of [26] propose a convex, tractable approach to determine the maximum power flexibility of an ADN by restricting the set of power exchange trajectories at the GCP to an ellipsoid, following (2) with $e$ the center and $E$ describing the axes and shape. Any point within this ellipsoid represents a vector of power exchanges at the ADN for the considered horizon. The use of (2) makes $\Omega_{p_{0}}$ explicit.
@@
In Section II-A we summarize important results from [26] before we extend them in the next sections to determine the flexibility potential for the provision of multiple AS. To this end, Section II-C shows the aggregation of multiple AS, Section II-E introduces cost-effectiveness, Section II-D proposes a probabilistic approach to integrate uncertainties and Section II-F extends the method for multiple feeders.
@@ II-A Network model and power aggregation
$p_{0}=G\begin{bmatrix}p\\ q\end{bmatrix}+b,\;|v|=[K_{p}K_{q}]\begin{bmatrix}p\\ q\end{bmatrix}+w,\;|i|=[J_{p}J_{q}]\begin{bmatrix}p\\ q\end{bmatrix}+d$ (3)
The coefficient matrix $G$ represents a linear mapping for the power balance, with constant coefficient $b$. Finally, $K_{p}$, $K_{q}$ and $J_{p}$, $J_{q}$ represent the linear coefficients mapping the power injections to the nodal voltage magnitudes $|v|$ and the branch current magnitudes $|i|$ respectively, with $w$ and $d$ the corresponding constant terms. Using the grid model (3), the network constraints are written as (4).
@@
$\max_{E,e,p}$ $\quad\textrm{log}(\textrm{det}(E)):\forall||\xi||\leq 1,\forall||\zeta_{t}||\leq 1,t=1..T$ (5a)
$\textrm{s.t.}\quad G$ $\begin{bmatrix}p\\ q\end{bmatrix}+B\zeta+b_{0}=E\xi+e,\;W\begin{bmatrix}p\\ q\end{bmatrix}\leq Z\zeta+z_{0}$ (5b)
$B$ and $b_{0}$ and $Z$ and $z_{0}$ are the affine mappings from $\zeta$ to $b$ and $\zeta$ to $z$ respectively. $W$ models the inequalities, including (4). Equality constraints should be eliminated when possible in robust optimization problems [35]. To this end, the authors of [26] exploit the structure of the problem. By writing $p=B_{1}x+B_{2}y$, with the columns of $B_{1}$ being an orthogonal basis for $G$ and the columns of $B_{2}$ spanning the null-space of $G$, equality constraints are eliminated. Finally, to make this problem tractable, an affine policy (6c) is introduced, with linear dependencies $K$ and $L_{t}$ on the position on the ellipsoid and the uncertainties respectively and a constant $\gamma$. With these reformulations and the introduction of $W_{1}=WB_{1}$, $W_{2}=WB_{2}$ $D=GB_{1}$, the problem is given by (6).
$\max_{E,e,K,\gamma,L_{t}}\quad\textrm{log}(\textrm{det}(E))$ (6a)
s.t. $W_{1}D^{-1}(E\xi+e-(B\zeta+b_{0}))+W_{2}y$
$\leq$ $Z\zeta+z_{0}\quad\forall||\xi||\leq 1,\forall||\zeta_{t}||\leq 1,t=1..T$ (6b)
$y=K\xi+\gamma+\sum_{t}L_{t}\zeta_{t}$ (6c)
To reformulate the robust constraints in (6) the worst case for each over the ellipsoidal aggregation and uncertainty sets is considered for each constraint individually. This results in the deterministic problem (7) where $\Theta_{t}$ represents the affine dependence on $\zeta_{t}$, $\nu$ collects the constant offsets, $w_{1,i}$, $w_{2,i}$ are the ith row of $W_{1}$, $W_{2}$ and $\alpha_{i}$ is an auxiliary variable.
$\max_{E,e,K,\gamma,L_{t}}\textrm{log}(\textrm{det}(E))$ (7a)
s.t. $\forall i=1..m\quad\alpha_{i}+w_{1,i}D^{-1}e+w_{2,i}\gamma-\nu_{i}\leq 0$ (7b)
@@ II-B Resource Constraints
The power flexibility aggregated at the GCP is provided by the controllable DERs in the ADN. Although their character is typically truly non-linear [36], resource capabilities may be accurately described through linear constraints, as shown for example in [37]. This allows a linear representation of all the constraints limiting the flexibility provision, as required to eliminate the equality constraints in (5). Here we only note the constraints for a BESS, but this can be generalized to other DERs. The particularity of BESSs is that they introduce time-coupling, meaning the flexibility aggregation problem can not be solved for each $t$ separately. The battery active and reactive power $p_{b},q_{b}$ and state of energy (SOE) constraints $\forall t$ are:
$p_{b}^{min}\leq$ $p_{b}\leq p_{b}^{max},\;SOE^{min}\leq SOE(t)\leq SOE^{max}$ (8a)
$q_{b}^{min}\leq$ $q_{b}\leq q_{b}^{max},\;SOE(t+1)=SOE(t)-p_{b}(t)\$(8b)

### v3

@@ I-B Flexibility aggregation approaches
We consider a DSO-TSO cooperation where the DSO aggregates the DER capabilities to offer flexibility to the TSO. The relevant problem is to quantify and explicitly determine the available flexibility, while accounting for network and resource constraints and stochastic boundary conditions.  [^1] In this respect, the considered problem is different than approaches presented in [45] and [13], where the operation of a distribution system under uncertainty is considered and works such as [30], where clearing of energy and flexibility markets is considered while accounting for DERs. In this section, existing approaches for flexibility aggregation are discussed and the remaining issues are identified. A fundamental distinction between the methods reported in literature can be made based on the type of algorithm used [14].
i) Simulation-based Approaches
The first group consists of approaches based on Monte Carlo (MC) simulations. Here, a significant number of scenarios are considered, each corresponding to a feasible actuation of the flexible resources. For each scenario, the corresponding power exchange at the GCP is obtained. If all constraints are satisfied, the corresponding power exchange is labeled as a feasible point within the PQ plane. By performing a large number of simulations (with a consequent computation time), an accurate mapping of the power capability can be made. An example of such an approach is presented in [20]. Recent developments (e.g. [40]) consider this problem and propose a Bayesian learning approach to alleviate the computational challenges. However, they still rely on an initial outer-box approximation of the flexibility set obtained through optimization.
@@
ii) Optimization Approaches: Exact Aggregation On the other hand, optimization-based approaches can be divided in a group of methods that attempt to compute the exact power aggregation set at the GCP and a group of methods that compute an approximate set, which can typically be described more easily. We further differentiate existing methods based on four important features of the flexibility aggregation sets: grid-awareness, time-coupling, cost-awareness and wether they account for uncertainty. As an example, [43] proposes an approach to compute the exact aggregated power flexibility set of multiple DERs without accounting for grid constraints. They also show that the number of constraints may be too large to solve the exact problem and have instead proposed a set of approximate models. Recognizing this issue, [46] proposes to cluster DERs based on the similarity of their capabilities and obtain an inner approximation through a specific type of polytope. The DERs in each cluster can then be exactly aggregated. Alternatively, in [44] a ”geometric prototype” is constructed by selecting a subset of the equations forming the exact power aggregation set, keeping geometrical features of the exact model while reducing the computational complexity. Approaches attempting to identify the exact aggregated flexibility set, while simultaneously satisfying the grid constraints, also exist. In this case, a set of OPF problems is typically solved. For instance, in [34] a set of non-approximated non-convex OPF problems are solved to obtain points on the aggregated capability curve in the PQ plane until a convergence criterion based on the distance between the obtained points is satisfied. This approach also allows to include cost constraints. Similar approaches are also proposed in [24], [8], [6] and [4]. Explicit computation of flexibility costs for these approaches is proposed in [33], but costs are computed a posteriori for identified points of the flexibility map and consider a single time step. These approaches only allow to compute the aggregation set for a single time step as the solved OPF problems cannot account for time coupling. To alleviate this problem, [23] proposes a network-informed data driven approach based on a classifier labeling samples to obtain an approximation allowing for temporal coupling.
iii) Optimization Approaches: Convex Approximations
@@
While the above methods can obtain flexibility areas closer to the exact set, they are only applicable when neglecting grid constraints or considering a single time step or have to resort to approximations to make the problem tractable. Therefore, methods seeking convex inner approximations of the exact flexibility set are desirable. An approach for the estimation of a robust aggregation set is presented in [36], where a linear power flow model is used. However, only single time steps are considered and DERs costs are not taken into account. In [7] a grid-aware approach, using a fixed-point linearization, is presented to compute the aggregate power flexibility over multiple time steps by approximating it as a hyperbox. However, uncertainty of stochastic prosumption is not accounted for. The authors attempt to model the costs, but only the one of the base trajectory, around which flexibility is offered, is included. In [9] the authors show that the flexibility aggregation set obtained through a hyperbox approximation is very conservative and that a larger one can be obtained with an inner ellipsoidal approximation. Using the same fixed-point linearization, the authors present tractable reformulations to obtain the maximum-volume ellipsoid under affine and quadratic disaggregation policies, while accounting for uncertainty through ellipsoidal uncertainty sets at each time step. An approximate approach decomposing the flexibility aggregation set in a ”virtual battery” and ”virtual generator” is presented in [41]. Uncertainty is accounted for through chance constraints assuming Gaussian probability distributions suitably transformed to deterministic constraints. The authors claim better results compared to the ellipsoidal inner approximation. The same authors extend this in [42] with a piecewise fitting of the aggregated cost function for the different time steps. However, the cost function is computed by minimizing the operational cost for different points on the time-decoupled aggregation sets and fitting an affine function. Therefore, it does not account for the time-coupling of the flexibility and its influence on the costs. An alternative approach to the approximate aggregation problem is proposed in [35]. By solving the disaggregation problem for samples selected in the flexibility aggregation space and labeling feasible points a convex ellipsoidal classifier is found. This ellipsoid is used as a surrogate solution space within which a polytope mapping the aggregated flexibility is determined. Uncertainty is accounted for when constructing the surrogate set.
I-C Contributions of the Paper
All the works above consider the flexibility aggregation problem to be the mapping of a single flexibility commodity at the ADN GCP. However, even though all flexibility services require power exchange, the provision of multiple ones (e.g. primary frequency regulation (FCR) and aFRR) have different implication on time-coupling and, therefore, the feasibility of an aggregation set. This is demonstrated in works regarding the optimal provision of AS using battery energy storage systems, such as [22]. Furthermore, in the works considering uncertainty, forecast errors are either modeled robustly, leading to overly conservative solutions, or through chance constraints related to individual time steps. This does not accurately represent the impact of stochastic variables. Finally, although some works consider costs, none of these methodologies enforce cost-effectiveness of flexibility in an appropriately time-coupled manner. To summarize, none of the above approaches consider simultaneously network constraints, time-coupling and cost of the flexibility and the impact of uncertainty. The provision of multiple AS is also not considered in state-of-the-art aggregation approaches. Therefore this paper makes the following contributions:
@@ I-C Contributions of the Paper
    Uncertainty is accounted for in a probabilistic way through joint uncertainty sets, following the P90 requirement [16] for the availability of power flexibility.
@@ II Problem Statement
The problem considered in this work is the flexibility aggregation at the GCP of DERs within ADNs with generic topologies (i.e. either meshed or radial) to provide multiple AS at the transmission level. To accurately model the available flexibility, it is essential to consider both the capabilities of DERs and the constraints of the network. This ensures the aggregated flexibility does not adversely impact the operation of the distribution system. Additionally, the impact of uncertain prosumption and the time-coupling of the DERs flexibility capabilities is highly relevant. Consider the general formulation (1), where $p_{0}$ is the vector representing the power exchanged at the GCP over all time steps, $\Omega_{p_{0}}$ is the set of feasible $p_{0}$ values, $p$ and $q$ are stacked vectors containing the vectors of controllable active and reactive power injections [^2] Note that in this formulation, the controllable injections are assumed to be known, with a capability curve that is constrained by the uncertainties. over all time steps and $\zeta\in\mathcal{U}$ is the stacked vector collecting the uncertainty drivers $\zeta_{t}$ over all time steps.
$\max_{\Omega_{p_{0}},p,q}$ $\quad J(\Omega_{p_{0}}):\forall\zeta\in\mathcal{U}$ (1a)
s.t. $G(p,q)+b(\zeta)=p_{0},\quad W(p,q)\leq z(\zeta)$ (1b)
$J$ represents the flexibility maximizing objective, $G$ and $b$ represent the mapping between respectively the controllable resources and the uncertainty drivers and the slack power, while $W$ and $z$ represent the network and resource constraints. The goal of (1) is to determine the flexibility-maximizing set $\Omega_{p_{0}}$ of power exchanges at the GCP. Any point within this set should be feasible for any realization of the uncertainty drivers. Directly obtaining such a set is in general untractable, even when neglecting grid constraints and using linear DER models [38]. The authors of [9] propose a convex, tractable approach to determine the maximum power flexibility of an ADN by restricting the set of power exchange trajectories at the GCP to an ellipsoid, following (2) with $e$ the center and $E$ describing the axes and shape. Any point within this ellipsoid represents a vector of power exchanges at the ADN for the considered horizon. The use of (2) makes $\Omega_{p_{0}}$ explicit.
@@
In Section II-A we summarize important results from [9] before we extend them in the next sections to determine the flexibility potential for the provision of multiple AS. To this end, Section II-C shows the aggregation of multiple AS, Section II-D proposes a probabilistic approach to integrate uncertainties, Section II-E introduces cost-effectiveness, and Section II-F extends the method for multiple feeders.
@@ II-A Network model and power aggregation
$p_{0}=G\begin{bmatrix}p\\ q\end{bmatrix}+b,\;|v|=K^{v}\begin{bmatrix}p\\ q\end{bmatrix}+w,\;|i|=K^{i}\begin{bmatrix}p\\ q\end{bmatrix}+d$ (3)
The coefficient matrix $G$ and constant coefficient $b$ represent an affine mapping for the power balance and $K^{v}$ and $K^{i}$ represent the linear coefficients mapping the power injections to the nodal voltage magnitudes $|v|$ and the branch current magnitudes $|i|$ respectively, with $w$ and $d$ the corresponding constant terms. Using the grid model (3), the network constraints are written as (4).
@@
[…]
$G$ $\begin{bmatrix}p\\ q\end{bmatrix}+M_{b}\zeta+b_{0}=E\xi+e,\;W\begin{bmatrix}p\\ q\end{bmatrix}\leq M_{z}\zeta+z_{0}$ (5b)
$M_{b}$ and $b_{0}$ and $M_{z}$ and $z_{0}$ are the affine mappings from $\zeta$ to $b$ and $\zeta$ to $z$ respectively. $W$ models the inequalities, including (4). Equality constraints should be eliminated when possible in robust optimization problems [18]. To this end, the authors of [9] propose to exploit the structure of the problem. By writing $p=B_{1}x+B_{2}y$, with the columns of $B_{1}$ being an orthogonal basis for $G$ and the columns of $B_{2}$ spanning the null-space of $G$, equality constraints are eliminated. Finally, to make this problem tractable, an affine policy (6c) is introduced, with linear dependencies $K$ and $L$ on the position on the ellipsoid and the uncertainties respectively and a constant $\gamma$. With these reformulations, the problem is given by (6).
[…]
$W_{1}D^{-1}$ $(E\xi+e-(M_{b}\zeta+b_{0}))+W_{2}y\leq M_{z}\zeta+z_{0}\quad$ (6b)
$y=K\xi+\gamma+L\zeta$ (6c)
$W_{1}=WB_{1},\;W_{2}=WB_{2},\;D=GB_{1}$ (6d)
To reformulate the robust constraints in (6) the worst case over the ellipsoidal aggregation set and uncertainty is considered for each constraint individually. Substituting (6c) in (6b) and evaluating for the worst case uncertainties gives an upper bound for the uncertain part (7c) denoted with the auxiliary variables $\alpha_{i}$ for each constraint $i$. Replacing this upper bound in (6b) gives the deterministic constraint (7b). Together, this results in the deterministic problem (7) where $\Theta$ represents the affine dependence on $\zeta$, $\nu$ collects the constant offsets and the subscript $i$ denotes the ith row.
$\max_{E,e,K,\gamma,L}\textrm{log}(\textrm{det}(E))\quad\textrm{s.t.}\quad\forall i=1..m$ (7a)
$\alpha_{i}+w_{1,i}D^{-1}e+w_{2,i}\gamma-\nu_{i}\leq 0$ (7b)
@@ II-B Resource Constraints
The power flexibility aggregated at the GCP is provided by the controllable DERs in the ADN. Although their character is typically truly non-linear [47], resource capabilities may be accurately described through linear constraints, as shown for example in [26]. This allows a linear representation of all the constraints limiting the flexibility provision, as required to eliminate the equality constraints in (5). As set of representative DERs, we consider BESSs, heat pumps and photovoltaic (PV) installations. The particularity of BESSs and HPs is that they introduce time-coupling, meaning the flexibility aggregation problem can not be solved for each $t$ separately. The battery active and reactive power $p_{b},q_{b}$ and state of energy (SOE) constraints $\forall t$ are:
$p_{b}^{min}\leq$ $p_{b}^{t}\leq p_{b}^{max},\;SOE^{min}\leq SOE^{t}\leq SOE^{max}$ (8a)
$q_{b}^{min}\leq$ $q_{b}^{t}\leq q_{b}^{max},\;SOE^{t+1}=SOE^{t}-p_{b}^{t}\Delta_{t}\$(8b)
The flexibility of HPs and PV installations is slightly more complex to model as the uncertainty of the heat demand and the global horizontal irradiance (GHI) directly impacts the capability of these DERs. The following linear model for HPs, coupled to a temperature-dependent heat demand with a heat buffer tank, is used. The variables $T$ and $Q$ represent the temperature and heat respectively, while subscripts $bt$, hphp, demanddemand respectively represent the buffer tank, heat pump and thermal demand. Finally, we define the buffer tank mass $m_{bt}$, the buffer tank losses $l_{bt}$, the heat pump model linear and constant terms $a_{hp}$ and $Q^{0}_{hp}$, the heat demand coefficient $h_{demand}$ and the comfort and environmental temperatures $T_{c}$ and $T_{\infty}$, related to the uncertainty vector by the affine mapping $H$.
$T_{bt}^{t+1}-T_{bt}^{t}=$ $\frac{\Delta t}{4200m_{bt}}(Q_{hp}^{t}-Q_{demand}^{t}-l_{bt})$ (9a)
$Q_{hp}^{t}=a_{hp}p_{hp}^{t}+$ $Q^{0}_{hp},\;p_{hp}^{min}\leq p_{hp}\leq p_{hp}^{max}$ (9b)
$Q_{demand}^{t}=h_{demand}$ $(T_{c}-T^{t}_{\infty}),\;T_{bt}^{min}\leq T_{bt}^{t}\leq T_{bt}^{max}$ (9c)
The uncertainty appears in the environmental temperature $T_{\infty}=T_{\infty}^{0}+H\zeta$. Similarly, the PV installations are modelled as follows, with the subscript $MPP$ denoting the maximum power point power and $M_{T}$ the affine mapping between the power and the uncertainty vector.
$0\leq p_{PV}^{t}\leq p_{PV}^{MPP,t},\;p_{PV}^{MPP,t}=p_{PV}^{MPP,t,0}+M_{PV}\zeta$ (10a)